Status Report

NASA Space Station On-Orbit Status 11 Nov 2003

By SpaceRef Editor
November 12, 2003
Filed under , , ,
NASA Space Station On-Orbit Status 11 Nov 2003

All ISS systems continue to function nominally, except as noted previously or below.

FE Alexander Kaleri continued the latest maintenance cycle on the Russian segment (RS) ventilation system, today cleaning the protective panel vent grilles of the FGB air ventilation system.  He also changed out the FGB’s two dust filters (PF1 & PF2).

CDR/SO Michael Foale conducted the first Increment 8 thirty-day functionality (“health”) test of the HRF GASMAP (Human Research Facility/Gas Analyzer System for Metabolic Analysis Physiology).  Last time done: 10/7.  [After unstowing, cabling and powering up the equipment, Ed let it run for 2 hrs. in standby mode to warm up, then performed a health check and reconfigured the system for a 2-3-hr. low power rundown prior to shutdown.  Afterwards, the GASMAP was deactivated, disconnected and stowed again.  The checkout used PuFF software along with GASMAP, in order to provide data for the ISS Environment Monitoring folks at MCC.]

Both crewmembers in turn took the periodic on-orbit hearing assessment (O-OHA) test, a NASA EHS (environmental health systems) examination to assess the efficacy of acoustic countermeasures. [The O-OHA test involves minimum audibility measurements for each ear over a wide range of frequencies and sound pressure levels, with the crewmembers using individual-specific Prophonics earphones, Bose ANC headsets and the SLM (sound level meter).  To conduct the testing, the experimenter is supported by special “EarQ” software on the MEC (medical equipment computer).  The baseline test is required for about Flight Day 14 for each new Expedition and is then performed once per month.  Last time done: 10/8.]

Mike Foale engaged in a teleconference with TsUP/Moscow specialists to discuss the performance of the LDI-11 laser range finder during the recent approach & docking of Soyuz TMA-3 on 10/20.  [Typical questions asked by the ground during Orbit 28400 were, What was the condition of the blister window in the Orbital Module during the flight and range-finding?  Was there any local fogging on the internal surface or between the panes?  How well could you see USOS lights in sunset?  How well could you see USOS lights and illuminated SM test area in the eclipse? Etc.]

The CDR conducted the regular weekly TVIS treadmill inspection, including its wire ropes.  The scheduled RED (resistive exercise device) inspection was postponed to 11/14 (tentatively), in order to accommodate a task concerning checkingdefibrillator battery voltage

In continuation of an activity started by Foale last week (11/6), he conducted a checkup of the defibrillator battery charge, by measuring open circuit voltage of battery #1007. [This measurement is required for the ground to verify that the required two “good” batteries for the defibrillator are in place after they normalizefrom charging.]

Both crewmembers completed their daily physical exercise program (2.5 hrs.) on TVIS treadmill, CEVIS cycle ergometer, and RED anaerobic exerciser.

TVIS exercise data files were downloaded by Alex Kaleri to the MEC (medical equipment computer), as its standard procedure whenever more than four sessions have been performed since last.

After Expedition 7 crewmembers provided some search locations, the Increment 8 crew was able to locate the “lost” RED canisters that are to be installed.  The canister swap has been rescheduled for tomorrow (11/12).

Foale transferred and rearranged resupply items for the HRF FOOT (Human Research Facility — foot/ground reaction forces during space flight) experiment from the E318 FOOT kit into other kits for ease of use during experiment operations.

Sasha Kaleri performed the regular weekly inspection of the BRPK air/condensate separator of the SRV-K water processing system, performed the daily routine maintenance of the SOZh life support system (including ASU toilet facilities) and also readied the IMS inventory system for the daily automated export/import of updates.

The Flight Engineer called down the “ad hoc” O2 partial pressure of the cabin air.  [O2 data for trending analyses by the ground are collected daily by the crew with the U.S. CSA-CP (compound specific analyzer-combustion products).]

Sasha also took weekly CO2 readings with the CDMK (carbon dioxide monitoring kit) and called down data and CDMK battery status.

Mike Foale set up the video equipment for recording today’s EPO (Educational Payload Operations) activity, which consisted of a demonstration of Newton’s Laws with EVA tools and flight features.  Video and voice of the block demo were downlinked in real time via Ku- and S-band for later use in educational products.

The CDR also conducted a checkout of the SPHERES BBT (Synchronized Position Hold, Engage, Reorient, Experimental Satellites — Beacon and Beacon Tester), including documentary recording with the Lab camcorder video setup.  For equipment checkout, Mike used the BBT to check the infrared and ultrasound environment.

Foale today worked on the hard drive repair for the SAMS (space acceleration measurement system), whose ICU (interim control unit) laptop did not properly boot up after it locked up on 10/24.  This activity had been postponed from the original scheduled date of 11/4 because of the longer-than-expected PMA-2 stowage transfers.  [Planned for today were partition repair on the laptop’s HDD (hard disk drive) and cleaning of the SAMS filter.  Today’s procedure was separated into two parts because the unattended portion of partition repair may take a while.]
FE Kaleri replace an emergency gas mask in the SM, swapping IPK-1 #3 with #4 from FGB stowage.  IPK-1 #3 was stowed and its location entered in the IMS.

TsUP/Moscow was today scheduled to conduct the regular periodic FGB solar array efficiency test (last time done: 4/3/03). [The periodic Russian efficiency testing keeps track of the energy-output performance of the photovoltaics over time under the degrading effects of the space environment (mostly from ultraviolet radiation and atomic oxygen).  Since the test requires the full power output of the solar arrays and the FGB itself does not have sufficient loads for drawing it, the U.S. side today increased loads on U.S. RACUs (Russian-to-American converter units) up to 1200 W, increasing and decreasing in steps of ~200 W each two minutes.]

The crew reported that three of seven stem adaptors identified in this weekend’s procedures for the Renal Stone experiment were broken. These adaptors attach to the UCDs (urine containment devices) to facilitate Renal Stone operations. The collection activities scheduled over the weekend were completed.  [This is not considered a threat to the activities during the rest of the Increment, as there are many other adaptors on board.] 
Yesterday’s PAO with veteran Skylab crewmembers was very successful.  The veteran astronauts, who were assembled at MSFC to commemorate the 30th anniversary of the launch of the last Skylab mission, asked many question of the ISS crew.

The PCS (portable computer system) laptop in the Joint Airlock (A/L) failed over the weekend.  The crew swapped out the hard drive in the A/L PCS with the last spare on board.  Additional on-board spare drives will be updated with current PCS software in the near term.

The Russia Services Group (RSG) has requested a one-hour time slot any day this week to perform testing of the replacement mission router PC at TsUP/Moscow. The primary router failed on 11/7, and the backup router was brought on line.The failed router has been replaced and needs to be tested to verify operability.  [When switching from one router to the other during the test period (i.e., from the new prime router to the backup routerto allow testing of the new one, thenback to the new prime after thetest), there will be an interruption of mission data services.]

Today’s CEO (crew earth observation) targets, limited in the current XPOP attitude by flight rule constraints on the use of the Lab nadir/science window, and including the targets of the Lewis & Clark 200-year memorial locations, were Industrialized SE Africa haze (looking right for views along the sub-continental escarpment, with different smog/smoke concentrations on the higher plateau versus the lower coastal plain), Lower Amazon River Basin (nadir pass over the Amazon estuary.  Looking left and right for detailed views), Mamore River, N Bolivia (Dynamic event.  Looking left into the sun-glint disc to document present positions of this major river on its floodplain.  Scattered evidence from ISS-7 shows significant changes in the positions of meanders occurred in less than 15 years [published on NASA’s “Earth Observatory” website on 14 June, 2003].  This unexpectedly fast rate of change suggests more thorough documentation of the Mamore River is called for. Glint views reveal high levels of detail), Patagonian glaciers (views of glacier tongues on both sides of the Andes are available on this pass over the northern ice field.  Smaller glacier tongues are requested.  Detailed handheld images have even revealed crevasse patterns.  Crevasse patterns indicate the style of ice flow in glaciers, most of which are believed to be moving more rapidly now than in the prior tens of thousands of years), and Patagonian Glaciers (second opportunity.  Views on the less cloudy, desert side of the Andes are available).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

SpaceRef staff editor.